# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v2

@@ 3.3.3 Performance Analysis of Fine-tuning-based Approaches
Figure 6: Loss curve and accuracy curve as the number of training epochs increases. We repeat the fine-tuning process three times and display the average loss and accuracy. For each repetition, we consistently used the same training and test sets, ensuring that any randomness in the results came solely from the fine-tuning process, not from variations in the training and test set splits. The final accuracies of the three replicates on the test set are 83.59%, 82.70%, and 83.93% respectively, and 83.40% on average. We add the best performance (red dash line) obtained using OpenAI’s embedding in §$\S$3.2 for comparison.
We report the loss curve and accuracy curve during the LoRA fine-tuning process in Figure 6. We observe an accuracy of 83.40% accuracy on the test data, which is marginally better than the best result (82.50% accuracy) using OpenAI embeddings in §$\S$3.2. However, such marginal improvements should not undermine the utility of fine-tuning. Comparing the fine-tuning of Llama-3-8b with the use of OpenAI embeddings is not straightforward due to inherent differences in models’ capability. To highlight the benefits of fine-tuning, we observe that the MLP classification model using embeddings derived from Llama-3-8b (79.40% accuracy in Table 6) performs worse than LoRA fine-tuned Llama-3-8b (83.40% accuracy). This demonstrates that fine-tuning Llama-3-8b can enhance performance compared to Llama-3-8b embeddings.
Similar to the case of the embedding-based classification model, we notice that the fine-tuning approach also suffers from some overfitting issues; see Figure 6. This is because Llama-3-8b is inherently complex; even though LoRA reduces the number of trainable parameters to 0.023% of the total parameters, the model remains complex. Common approaches to address over-fitting include reducing model complexity, adding regularization, and increasing the dataset size. We experimented with various LoRA configurations, such as further reducing the number of trainable parameters, increasing the dropout ratio, and applying larger weight decay in an attempt to mitigate over-fitting. However, these modifications generally reduced performance on the training set without yielding significant improvements on the test set.
We also consider two other versions of the LoRA fine-tuned model to document their performance:
  • $\bullet$
    Larger dataset: One way to improve performance and alleviate overfitting is to fine-tune the model on a larger dataset that also includes data on headline pairs that are not significantly different from each other. In our setting, including insignificant pairs of headlines expands the dataset from 39,158 pairs to 134,941 pairs. However, even with this expanded training data, accuracy on the test set continues to hover around 83%, which indicates that incorporating more training data does not necessarily lead to better performance, at least for the current dataset. We refer interested readers to Web Appendix §$\S$§D.2.1 for details of this analysis and numerical results.
  • $\bullet$
    DistilBERT Fine-tuning: Before the widespread adoption of the term “LLM”, BERT was considered one of the best language models available, and it is still widely used as a benchmark in many NLP tasks. To provide a comparison with these earlier models, we therefore fine-tune DistilBERT – a distilled version of BERT that retains over 95% of BERT’s performance while containing 40% fewer parameters (sanh2019distilbert; devlin2018bert). Fine-tuned DistilBERT provides an accuracy of about 75% on the test dataset, which is good, but still lower than that of the fine-tuned Llama-3-8b model. Please see Web Appendix §$\S$§D.2.2 for details on the implementation and performance of the fine-tuned DistilBERT.
In sum, we find that using more data and/or different base models does not lead to significant improvement in performance in our setting.
@@ 3.4 Summary of Performance of Pure-LLM Based Methods
We summarize the performance of various pure-LLM-based methods in Table 7. We also compare the accuracy of human respondents on the same task. Details of the survey questions and the accuracy analysis of human respondents are shown in Web Appendix §$\S$§A.
Method Accuracy Notes
Human respondents 51.58% On 4,571 human labeled pairs
Prompt-based approach $<$65% GPT-4 with in-context learning, evaluated on 1,000 random samples
Classification model with LLM text embeddings 82.50% OpenAI-3072E embedding model with logistic regression
Fine-tuned LLM model 83.40% LoRA fine-tuned Llama-3-8b
@@
Table 7: Summarization of performance of various pure-LLM-based methods and comparison with survey results. All results in this table have been reported before. We only report the best performance of each method.
@@
There are four main takeaways from these comparisons. First, human respondents cannot easily predict which headline would be more engaging and likely to lead to higher CTRs. This is consistent with the fact that media firms (and even expert editors) rely on experimentation to decide which headlines and articles to display to users. Second, while simple prompt-based approaches like in-context learning with examples are marginally better than purely random guesses or human predictions, they nevertheless do not provide sufficiently high accuracy. Third, both embedding-based models and fine-tuning perform equally well (at least in our setting), and provide equally good performance with around 83% accuracy. They each have their pros and cons. The main advantage of the logit classification model with OpenAI embeddings is that it is extremely easy to implement and use in business settings since it does not require any significant model training/deployment. However, embeddings from the OpenAI models are proprietary, and as such, we lack visibility into the exact nature of the embeddings. In contrast, fine-tuning approaches use open-source models such as Llama-3-8b, and we have access to the underlying parameters of the LLM. However, this transparency comes with significant costs – fine-tuning LLMs can be computationally expensive and requires non-trivial training and deployment skills that not all businesses may have or want to invest in. The fourth and most important takeaway is that while pure-LLM-based approaches are informative, they are unable to reach the accuracy of experimental approaches. That is, firms cannot replace content experimentation with LLM-based predictions.
@@
Finally, we discuss the monetary costs associated with three LLM-based approaches. The cost of prompt-based, embedding-based, and LoRA fine-tuning in our analysis are around \$20, \$0.18, \$3, respectively. The cost of prompt-based and embedding-based comes from the access to OpenAI’s API in order to get responses and text embeddings. The cost of fine-tuning comes from renting a cloud computing server equipped with one NVIDIA A100 GPU. Overall, we find that the prompt-based approaches are the most expensive, while embedding and fine-tuning are the cheapest in terms of cost. However, in terms of run time and effort, fine-tuning is more costly. The detailed comparison of the cost, run-time, and engineering effort required for each pure-LLM method are discussed in detail in Web Appendix §$\S$§E.
@@ 4 Our Approach: LLM-Assisted Online Learning Algorithm
So far, we have seen that pure-LLM-based approaches can help with headline selection with reasonably high accuracy (around 83% for the embedding-based and LoRA fine-tuning approaches). The main advantage of pure LLM-based approaches is that they can be deployed without any experimentation, i.e., the manager can simply use the LLM-based model to predict which headline is better and go with it. Thus, there is no need to use any traffic for experimentation. The downside, of course, is that these models are not fully accurate – in a pairwise comparison, in around 17% of the cases, they are likely to mispredict which headline is better. This can, of course, lead to lower clicks and revenues for the firm. [^15] It should be noted that even this accuracy is calculated under the assumption that the results from the A/B tests are the ground truth. However, even if we were to reach 100% accuracy on this dataset, we still may not reach 100% accuracy in practice because of Type I errors in the creation of significant samples. Thus, it is not clear that relying on pure-LLM-based methods will lead to better overall outcomes for the firm.
@@ 4.2 LOLA
In the first step, we use LLMs to get an initial prediction of each headline’s CTR. Note the prediction task here is somewhat different from that in §$\S$3. In the previous section, we used LLM-based approaches to perform pairwise comparisons of headlines and predict which headline amongst a pair is more engaging. However, as discussed earlier, that approach does not work when we have more than two headlines (which is the case for over 90% of the A/B tests in our data). Further, we need actual CTR predictions for each headline to feed into the second step of our algorithm. Therefore, we adapt the LLM-based classification models from §$\S$3 for predicting CTRs by making the following modifications: changing the loss function from log loss to mean squared error (MSE), using single headline inputs instead of headline pairs, and switching from binary output to continuous CTR prediction. The training and fine-tuning procedures remain unchanged. [^16] Using CTR as the outcome measure does not adversely affect the performance. For instance, using linear regression with 256-dimensional text embeddings from OpenAI to predict CTR for significant pairs and labeling the headline with the higher predicted CTR as the winner yields an accuracy of around 75%, which is very close to the 77.31% accuracy obtained with logistic regression. This minor difference is likely due to the inherently skewed nature of the CTR distribution and the sensitivity of MSE to outliers.
We compare two approaches for CTR prediction: an OpenAI-embedding-based method and a LoRA fine-tuned Llama-3-8b model. Our numerical results indicate that the R-squared values for the embedding-based approach and the LoRA fine-tuned method are 0.241 and 0.315, respectively. When selecting the headlines with the highest predicted CTRs, the accuracy—defined as the percentage of times the chosen headline indeed has the highest actual CTR—is 42.66% for the embedding-based method and 46.55% for the LoRA fine-tuned method. Note that this accuracy is significantly lower than the binary classification accuracy discussed in §$\S$3 because we can have more than two headlines in a test and many headlines have no significant difference in CTRs. Based on these results, we select the LoRA fine-tuned Llama-3-8b model as the best CTR prediction model in the first step. Additional details on the CTR prediction methods and their results can be found in Web Appendix §$\S$§F.
@@
The 2-UCBs algorithm offers three advantages. First, it is easy to implement, requiring the tuning of only two hyperparameters. Second, it can be used in conjunction with any LLM-based CTR prediction model in the first step and any outcome of interest in the second step. We use clicks as the outcome of interest given our application setting and data; however, other outcomes, such as time spent on a page or other measures of engagement, can be easily used instead of clicks. Finally, it is intuitive and easy to implement. Because it can easily accommodate the edge cases, e.g., one where the manager has no auxiliary information or one where the manager wants to fully rely on the LLM predictions, there is no need for extra engineering effort to set up separate systems for different edge cases. This adaptability is extremely valuable given the rapid advances in LLM technology. When the capability of the pure-LLM method continues to advance, the manager can easily fine-tune and increase $n^{\text{aux}}$ to take more advantage of LLM.
@@ 4.3 Benchmarks and Implementation Details
    Pure LLM-based approach: In this approach, the firm uses LLMs to predict the CTR for all the arms (headlines) and then allocates all the traffic to the arm with the highest predicted CTR. This is equivalent to a fully greedy policy based on LLM predictions. We use the training data to learn a model that predicts the CTR of a headline based on an LLM (either through text embeddings or LoRA fine-tuning). In our numerical study, we use the best-performing pure-LLM-based approach based on our experiments (see Web Appendix §$\S$§F), i.e., the LoRA fine-tuned Llama-3-8b with MSE loss, which gives a 46.55% accuracy. [^17] Note that since this is a pure prediction-based greedy approach, there are no hyper-parameters associated with the algorithm, and hence we do not use the fine-tuning data.
@@
    LOLA: This is our proposed approach. In the first step, we use CTR predictions from the best-performing pure-LLM-based CTR prediction model, i.e., the LoRA fine-tuning Llama-3-8b with MSE loss. Note that this is the same model that we use for the pure LLM-based approach, which allows for a fair comparison of the two approaches. For the second step, we use the LLM-2UCBs algorithm as outlined in §$\S$4.2. We fine-tune the combination of the two hyperparameters on the fine-tuning data, and we consider $n^{\text{aux}}\in\{600,800,1000,1200,1400\}$ and $\alpha\in\{0.02,0.04,\dots,0.10\}$. Based on fine-tuning, we choose $n^{\text{aux}}=1000$ and $\alpha=0.08$ for our setting.
    Note that the former two methods can be viewed as special cases of our proposed LOLA algorithm. Both our algorithm and the pure LLM-based approach use the exact same CTR prediction algorithm. Hence, the pure LLM-based approach (a.k.a pure exploitation, greedy algorithm) is a special case of LOLA, wherein we set $n^{\text{aux}}=\infty$, indicating complete trust in the LLM prediction results. This method fully relies on the LLM CTR prediction, directly selecting the headline with the highest predicted CTR throughout the entire horizon. Similarly, the standard UCB algorithm is a special case of LOLA with $n^{\text{aux}}=0$, indicating no reliance on LLM CTR predictions. Again, note that we use the same hyper-parameter $\alpha=0.08$ across both algorithms, which ensures that our algorithm defaults to the standard UCB with $\alpha=0.08$ when we ignore the CTR predictions from the LLM.
@@ 4.4 Numerical Results
Figure 7: Average clicks per experiment per period under different time horizon multipliers. Note that the Y-axis captures the average clicks per test per period. For instance, if there is a test with two headlines conducted under $x=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$. The Y value is simply the average of this number $0.03$ over all tests. This measure scales well with the platform’s total clicks in tests because headlines in different tests with different numbers of headlines take the same weight in this measure.
We visualize the results from our numerical simulations in Figure 7 and document the pairwise comparison of different algorithms, along with the relative percentage reward improvement and the p-values from the t-tests in Table 8. We find that, on average, LOLA performs the best among all algorithms across all time horizons. We document a few additional patterns of interest.
@@
Table 8: Percentage % reward improvement between two algorithms with significance levels: $p\leq$ 0.05, ∗∗ $p\leq$ 0.01, ∗∗∗ $p\leq$ 0.001, ∗∗∗∗ $p\leq$ 0.0001. For example, under the impression per headline equal to 50, the average click per test of LOLA and UCB is 0.053760 and 0.050267, respectively; so the relative improvement is calculated as ${0.053760}/{0.050267}-1=6.95\%$. Parameter $\tau$ represents the average impression per headline to scale the time horizon $T=\tau\times K$.
@@
First, the Pure LLM approach exhibits consistent performance across different $\tau$ values because it is not an adaptive algorithm. The small fluctuation at $\tau\in\{50,100,200\}$ is due to the randomness of average clicks per period, which tends to be larger under the smaller time horizon. For the other three approaches, which actively learn the CTR during the experiments, their average performance per period increases as the time horizon extends. This improvement is because the number of arms in the experiment remains constant, and with increasing periods, these algorithms have more opportunities to learn the CTRs, thus achieving better performance. Comparing the performance of LOLA with pure LLM methods, we find that LOLA generally outperforms Pure LLM, except when $\tau=50$, where there is no significant difference between LOLA and Pure LLM. This is because the short planning horizon does not allow LOLA to learn new information over and above the information already available from the LLM. As $\tau$ increases to 1000, we see that LOLA significantly outperforms Pure LLM, generating 5.02% more clicks. This suggests that if the horizon is very short, the planner/firm will not see significant improvement over simply relying on LLM predictions. However, as the horizon increases, LOLA will start showing significant gains.
@@
Next, comparing LOLA with regular UCB, we see that it always outperforms UCB. Interestingly, here we see that, as the time horizon increases, the relative improvement from LOLA diminishes. Intuitively, this is because both UCB and LOLA are active learning algorithms, and as the time horizon increases, they are both able to learn the CTRs better. While LOLA starts with a relative information advantage, as the time horizon increases, this advantage diminishes. Formally, this pattern is due to the asymptotic optimality of the UCB and LOLA algorithms, implying that as the time horizon approaches infinity, the initialization has a negligible effect on the average performance over an infinitely long period. We also see that E&C consistently underperforms compared to LOLA because it neither leverages LLM predictions nor achieves asymptotic optimality. While E&C shows better performance as the horizon increases, even when $\tau=1000$, LOLA is still better than E&C by 5.08%, which is a large improvement in the media and publishing industry.
@@
Finally, comparing the benchmark algorithms with each other, we see that when the horizon is short, the Pure LLM does better, while UCB is better when the horizon is longer. In general, this reflects the intuition that active learning is not very useful when the time horizons are very short since the algorithm has no time to explore and learn the CTRs effectively (and then exploit this learning). So simply going with the pure LLM predictions is better. In contrast, when the horizons are longer, UCB (or active learning) outperforms using static predictions from the LLM. Note that LOLA combines the strengths of both UCB and Pure LLM, and hence it is able to outperform both and provide superior performance in both short and long horizons. We also note that E&C is generally worse than both Pure LLM and UCB, which suggests that the current Upworthy practice is suboptimal. This is because E&C neither exploits LLM information and nor does it actively learn over time.
@@
In summary, we find that LOLA outperforms existing approaches for experimentation by leveraging the strengths of LLMs and combining them with the benefits of active/online learning. In particular, our results demonstrate the value of LOLA as the time horizon increases, and highlight the limitations of relying solely on pure LLM predictions or the naive E&C approach.

### v3

@@ 3.3.1 Performance Analysis of Fine-tuning-based Approaches
Figure 4: Loss curve and accuracy curve as the number of training epochs increases.
We report the loss curve and accuracy curve during the LoRA fine-tuning process in Figure 4. We observe an accuracy of 46.86% on the test data, which is marginally better than the best performance of 46.28% using OpenAI embeddings in §$\S$3.2. However, such similar performance should not undermine the utility of fine-tuning. Comparing the fine-tuning of Llama-3-8b with the use of OpenAI embeddings is not straightforward due to inherent differences in the models’ capabilities. To highlight the benefits of fine-tuning, we observe that the Linear CTR prediction model using embeddings derived from Llama-3-8b (42.78% on the test data, reported in Web Appendix §$\S$§C.2 Table A2) performs worse than the LoRA fine-tuned Llama-3-8b. This demonstrates that fine-tuning Llama-3-8b can enhance performance compared to using Llama-3-8b embeddings alone.
In Figure 4, we observe that the accuracy on the test set (46.86%) is higher than on the training set (42.09%), which is uncommon in machine learning tasks. The reason for this discrepancy might be that the split of the training and test datasets results in some easily distinguishable headlines being included in the test set, making the training set more challenging than the test set. To verify our hypothesis, in Web Appendix §$\S$§D.4, we try several additional random splits of the training and test datasets and evaluate their performance on them. We find that in these other splits, the test performance is slightly worse than the training performance.
@@ 3.4 Summary of Performance of Pure-LLM Based Methods
We summarize the performance of various pure-LLM-based methods in Table 6. We also report the results of human respondents on a similar task but with paired headline comparison. Details of the survey questions and the accuracy analysis of human respondents are shown in Web Appendix §$\S$§A.
Method Accuracy on test set Notes
Random guess 33.02%
Human respondents - Similar to random guess as tested on 4,571 human responses
Prompt 39.96% GPT-4 with in-context learning
Embedding 46.28% OpenAI-256E embedding model with linear regression
Fine-tuning 46.86% LoRA fine-tuned Llama-3-8b
@@
Table 6: Summarization of performance of various pure-LLM-based methods. All results in this table have been reported before. We only report the best performance of each method.
@@
There are four main takeaways from these comparisons. First, human respondents cannot easily predict which headline would be more engaging and likely to lead to higher CTRs. This is consistent with the fact that media firms (and even expert editors) rely on experimentation to decide which headlines and articles to display to users. Second, while simple prompt-based approaches like in-context learning with examples are better than purely random guesses or human predictions, they nevertheless do not provide sufficiently high accuracy. Third, both embedding-based models and fine-tuning perform equally well (at least in our setting), and provide equally good performance with around 46-47% accuracy on the test set. They each have their pros and cons. The main advantage of the linear CTR prediction model with OpenAI embeddings is that it is easy to implement and use in business settings since it does not require any significant model training/deployment. However, embeddings from the OpenAI models are proprietary, and as such, we lack visibility into the exact nature of the embeddings. In contrast, fine-tuning approaches use open-source models such as Llama-3-8b, and we have access to the underlying parameters of the LLM. However, this transparency comes with costs – fine-tuning LLMs requires non-trivial training and deployment skills that not all businesses may have or want to invest in. The fourth and most important takeaway is that while pure-LLM-based approaches are informative, they are unable to reach the accuracy of experimental approaches. That is, firms cannot replace content experimentation with LLM-based predictions.
An important caveat here is consumer preferences may have changed significantly since the time period of the experiment (2013–2015). As such, the prediction numbers shown here may not represent LLMs’ ability to predict the preferences of consumers today. [^11] This concern is somewhat assuaged by Web Appendix §$\S$§C.4, which shows that consumer preferences did not change significantly during the 2013–2015 timeframe (at least within the prediction capabilities of the LLMs). However, we cannot empirically comment on the period beyond 2015. As such, the performance of LLM-based methods in §$\S$3 should be mainly used for relative comparisons with each other. More broadly, both LLM capabilities and consumer preferences will continue to evolve. Therefore, if a firm/manager is concerned about the generalizability/accuracy of a specific LLM-based prediction model to their current business environment, the simplest solution is to retrain their LLM-based models more frequently and collect up-to-date labeled data to maintain relevance.
@@
Finally, we discuss the monetary costs associated with three LLM-based approaches. The cost of prompt-based, embedding-based, and LoRA fine-tuning in our analysis are around \$15.87, \$0.29, \$3, respectively. The cost of prompt-based and embedding-based comes from the access to OpenAI’s API in order to get responses and text embeddings. The cost of fine-tuning comes from renting a cloud computing server equipped with one NVIDIA A100 GPU. Overall, we find that the prompt-based approaches are the most expensive, while embedding and fine-tuning are the cheapest in terms of cost. However, in terms of run time and effort, fine-tuning is more costly. The detailed comparison of the cost, run-time, and engineering effort required for each pure-LLM method are discussed in detail in Web Appendix §$\S$§E.
@@ 4 Our Approach: LLM-Assisted Online Learning Algorithm
So far, we have seen that pure-LLM-based approaches can help identify the best headline with 46.86% accuracy using the LoRA fine-tuning approach. The main advantage of pure LLM-based approaches is that they can be deployed without any experimentation, i.e., the manager can simply use the LLM-based model to predict which headline is the best and go with it. Thus, there is no need to use any traffic for experimentation. The downside, of course, is that these models are not fully accurate – in some cases, they are likely to mispredict which headline is the best. This can, of course, lead to lower clicks and revenues for the firm.Thus, it is not clear that relying on pure-LLM-based methods will lead to better overall outcomes for the firm.
@@ 4.2 LOLA
In the first step, we use LLMs to get an initial prediction of each headline’s CTR. Based on numerical results in §$\S$3, we select the best-performing approach, LoRA fine-tuned Llama-3-8b model, as the best CTR prediction model in the first step.
@@
The intuition behind using two UCBs instead of a single UCB incorporating LLM predictions $U_{k}^{2}$ is as follows: if we rely solely on $U_{k}^{2}$, and the LLM’s CTR predictions are significantly incorrect, i.e., overestimating the CTR of a poorly performing headline (underestimating the CTR of the best headline is of less risk because there are more suboptimal headlines than the one and only best headline), it would take many rounds to correct this error, especially when the number of auxiliary samples, $n^{\text{aux}}$ is large. On the other hand, the first UCB $U_{k}^{1}$, though unbiased, suffers from large stochastic noise, requiring large $n_{k}^{t}$ to get an accurate estimator. By choosing the minimum of the two UCBs, we balance these opposing risks effectively. However, the formalism for this approach is only partially addressed in prior literature (gur2022adaptive), which operates under a different setting without LLMs. In particular, (gur2022adaptive) shows 2-UCBs is optimal under the assumption that the auxiliary data is generated from a linear function class, including the standard UCB setting with exogenous auxiliary samples. The formalism in our context would require additional assumptions about the errors in the LLM’s predictions, which remains an open research question with limited understanding. We leave this formalization for future work.
The LOLA algorithm offers three advantages. First, it is easy to implement, requiring the tuning of only two hyperparameters. Second, it can be used in conjunction with any LLM-based CTR prediction model in the first step and any outcome of interest in the second step. We use clicks as the outcome of interest given our application setting and data; however, other outcomes, such as time spent on a page or other measures of engagement, can be easily used instead of clicks. Finally, it is intuitive and easy to implement. Because it can easily accommodate the edge cases, e.g., one where the manager has no auxiliary information or one where the manager wants to fully rely on the LLM predictions, there is no need for extra engineering effort to set up separate systems for different edge cases. This adaptability is valuable given the rapid advances in LLM technology. When the capability of the pure-LLM method continues to advance, the manager can easily fine-tune and increase $n^{\text{aux}}$ to take more advantage of LLM.
As a final note, our algorithm does not update the LLM during the online learning phase. Instead, we use the LLM fine-tuned on historical data solely for initialization purposes. That being said, the algorithm can be readily modified to fine-tune LLMs with online data collected during the learning phase. While such fine-tuning could potentially improve LOLA’s performance, the improvement is expected to be limited in our current dataset. As shown in Figure A3 in Appendix §C.3, achieving a significant performance boost in LLMs would require a much larger dataset than is currently available to us.
Discussion: We now provide a broader discussion on LOLA. We note that the general idea that priors can impact regret and rewards is not new and has existed for some time; see bubeck2013prior and russo2014learning. These earlier papers primarily focus on theoretical analyses and often rely on assumptions about the quality of priors. However, since we do not have a theoretical understanding of LLM predictions, it is unclear how these predictions can be translated to priors or the nature of such priors. Separately, the 2-UCBs algorithm by gur2022adaptive was designed for a setting where the experimenter has access to auxiliary data. However, their algorithm assumes that the researcher has access to both the size of this auxiliary sample ($n^{aux}$) and knowledge of the prediction error in this sample. Thus, neither of these earlier approaches is directly applicable to our setting.
A key challenge in our setting is that the LLM predictions cannot be naively treated as priors since we do not know their prediction error or theoretical properties. The main novelty of LOLA comes from the idea that we can treat the LLM predictions as coming from a “pseudo-sample” and that we can fine-tune the size of this pseudo-sample ($n^{aux}$) for a given application based on prior data. Thus, we build on the 2UCBs algorithm but modify it to accommodate LLM predictions – that do not come from a real sample and whose quality and theoretical properties are unknown.
@@ 4.3 Benchmarks and Implementation Details
    Pure LLM-based approach: In this approach, the firm uses LLMs to predict the CTR for all the arms (headlines) and then allocates all the traffic to the arm with the highest predicted CTR. This is equivalent to a fully greedy policy based on LLM predictions. We use the training data to learn a model that predicts the CTR of a headline based on an LLM (either through text embeddings or LoRA fine-tuning). In our numerical study, we use the best-performing pure-LLM-based approach based on our experiments, i.e., the LoRA fine-tuned Llama-3-8b, which gives a 46.86% accuracy. [^12] Note that since this is a pure prediction-based greedy approach, there are no hyper-parameters associated with the algorithm, and hence we do not use the fine-tuning data.
@@
    LOLA: This is our proposed approach. In the first step, we use CTR predictions from the best-performing pure-LLM-based CTR prediction model, i.e., the LoRA fine-tuning Llama-3-8b with MSE loss. Note that this is the same model that we use for the pure LLM-based approach, which allows for a fair comparison of the two approaches. For the second step, we use the LLM-2UCBs algorithm as outlined in §$\S$4.2. We fine-tune the combination of the two hyperparameters on the fine-tuning data, and we consider $n^{\text{aux}}\in\{600,800,1000,1200,1400\}$ and $\alpha\in\{0.02,0.04,\dots,0.10\}$. Based on fine-tuning, we choose $n^{\text{aux}}=1000$ and $\alpha=0.08$ for our setting. We refer interested readers to Web Appendix §$\S$§F for a detailed sensitivity analysis of $n^{\text{aux}}$.
    Note that the first two methods can be viewed as special cases of our proposed LOLA algorithm. Both our algorithm and the pure LLM-based approach use the exact same CTR prediction algorithm. Hence, the pure LLM-based approach (a.k.a pure exploitation, greedy algorithm) is a special case of LOLA, wherein we set $n^{\text{aux}}=\infty$, indicating complete trust in the LLM prediction results. This method fully relies on the LLM CTR prediction, directly selecting the headline with the highest predicted CTR throughout the entire horizon. Similarly, the standard UCB algorithm is a special case of LOLA with $n^{\text{aux}}=0$, indicating no reliance on LLM CTR predictions. Again, note that we use the same hyper-parameter $\alpha=0.08$ across both algorithms, which ensures that our algorithm defaults to the standard UCB with $\alpha=0.08$ when we ignore the CTR predictions from the LLM.
  • $\bullet$
    UCB with LLM priors: To demonstrate the effectiveness of the 2-UCBs policy $\min\{U^{1},U^{2}\}$ compared to using only the second UCB $U^{2}$, we also run the UCB with LLM as priors, i.e., play the arm $a_{t}=\arg\max_{k\in[K]}U_{k}^{2}$. This comparison highlights how selecting the minimum of the two UCBs improves performance by mitigating the risk of overestimated CTRs from the LLM predictions.
@@ 4.4 Numerical Results
Figure 5: Average clicks per experiment per period under different time horizon multipliers. Note that the Y-axis captures the average clicks per test per period. For instance, if there is a test with two headlines receiving 1 and 2 clicks, respectively, under $\tau=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$. The Y value is simply the average of this number $0.03$ over all tests. This measure scales well with the platform’s total clicks in tests because headlines in different tests with different numbers of headlines take the same weight in this measure.
We visualize the results from our numerical simulations in Figure 5131313The Y metric in Figure 5 is essentially the accumulated reward divided by $\tau$, i.e., the value of the X-axis. We choose to report this Y metric instead of the actual regret for two reasons: (1) the ground truth arm is not known in practice, making it impossible to compute regret; and (2) for better visualization. Specifically, the accumulated reward (unscaled by $\tau$) grows very rapidly as $\tau$ increases, which compresses the visual differences between algorithms, making them difficult to distinguish with the naked eye. With this being said, the results in Table 7 can still be interpreted as the percentage improvement of the accumulated reward because all algorithms are measured at the fixed $\tau$ and the accumulated reward $=$ Y ×$\times\tau$., and document the pairwise comparison of different algorithms, along with the relative percentage reward improvement and the p-values from the t-tests in Table 7. We find that, on average, LOLA performs the best among all algorithms across all time horizons.141414 Note that we cannot translate the click improvement to revenues or other business outcomes of interest since we do not have data on how clicks/time spent on the website relate to downstream metrics like advertising and subscription revenues. Therefore, we do not make additional claims on the business impact of these numbers. That said, if the firm has access to these numbers, it should be relatively straightforward to make that connection.
@@
Table 7: Percentage improvement in reward (total clicks) between the algorithms LOLA (LLM-2UCBs), UCB, Pure LLM, and E&C (we omit UCB with LLM priors here due to space constraints, but report its performance relative to LOLA in Figure 5). Significance levels: $p\leq$ 0.05, ∗∗ $p\leq$ 0.01, ∗∗∗ $p\leq$ 0.001, ∗∗∗∗ $p\leq$ 0.0001. For example, under the impression per headline equal to 50, the average click per test of LOLA and UCB is 0.053760 and 0.050267, respectively; so the relative improvement is calculated as ${0.053760}/{0.050267}-1=6.95\%$. Parameter $\tau$ represents the average impression per headline to scale the time horizon $T=\tau\times K$.
@@
We now document a few additional patterns of interest. First, the Pure LLM approach exhibits consistent performance across different $\tau$ values because it is not an adaptive algorithm. The small fluctuation at $\tau\in\{50,100,200\}$ is due to the randomness of average clicks per period, which tends to be larger under the smaller time horizon. For the other three approaches, which actively learn the CTR during the experiments, their average performance per period increases as the time horizon extends. This improvement is because the number of arms in the experiment remains constant, and with increasing periods, these algorithms have more opportunities to learn the CTRs, thus achieving better performance. Comparing the performance of LOLA with pure LLM methods, we find that LOLA generally outperforms Pure LLM, except when $\tau=50$, where there is no significant difference between LOLA and Pure LLM. This is because the short planning horizon does not allow LOLA to learn new information over and above the information already available from the LLM. As $\tau$ increases to 1000, we see that LOLA significantly outperforms Pure LLM, generating 5.02% more clicks. This suggests that if the horizon is very short, the planner/firm will not see significant improvement over simply relying on LLM predictions. However, as the horizon increases, LOLA will start showing significant gains.
@@
Second, comparing LOLA with regular UCB, we see that it always outperforms UCB. Interestingly, here we see that, as the time horizon increases, the relative improvement from LOLA diminishes. Intuitively, this is because both UCB and LOLA are active learning algorithms, and as the time horizon increases, they are both able to learn the CTRs better. While LOLA starts with a relative information advantage, as the time horizon increases, this advantage diminishes. Formally, this pattern is due to the asymptotic optimality of the UCB and LOLA algorithms, implying that as the time horizon approaches infinity, the initialization has a negligible effect on the average performance over an infinitely long period. We also see that E&C consistently underperforms compared to LOLA because it neither leverages LLM predictions nor achieves asymptotic optimality. While E&C shows better performance as the horizon increases, even when $\tau=1000$, LOLA is still better than E&C by 5.08%, which is a large improvement in the media and publishing industry.
@@
Third, comparing the benchmark algorithms with each other, we see that when the horizon is short, the Pure LLM does better, while UCB is better when the horizon is longer. In general, this reflects the intuition that active learning is not very useful when the time horizons are very short since the algorithm has no time to explore and learn the CTRs effectively (and then exploit this learning). So simply going with the pure LLM predictions is better. In contrast, when the horizons are longer, UCB (or active learning) outperforms using static predictions from the LLM. Note that LOLA combines the strengths of both UCB and Pure LLM, and hence it is able to outperform both and provide superior performance in both short and long horizons. We also note that E&C is generally worse than both Pure LLM and UCB, which suggests that the current Upworthy practice is suboptimal. This is because E&C neither exploits LLM information and nor does it actively learn over time.
Finally, we will discuss the performance of UCB with LLM priors. Comparing LOLA (2-UCBs) with UCB with LLM priors, the key difference lies in whether the regular UCB $U^{1}$ is included in the algorithm. The core trade-off here is between the large potential error of $U^{2}$ when the LLM’s predictions are incorrect and the stochastic noise in $U^{1}$, especially in the early stages. Our observations confirm this trade-off. In short time horizons $\tau=50$, LOLA underperforms compared to UCB with LLM priors, by a significant margin of $-0.74\%$. This is because the error in the LLM-based CTR predictions is smaller than the stochastic noise of the regular CTR estimator at the start, and including the noisier $U^{1}$ reduces performance. However, as the time horizon increases (e.g., $\tau=800$ and $\tau=1000$), LOLA significantly outperforms UCB with LLM priors by $0.37\%$ and $0.47\%$, respectively. This improvement occurs because, over longer periods, the error in LLM-based CTR predictions becomes more significant than the stochastic noise of the regular estimator, which can also be observed by comparing the performance between the standard UCB and Pure LLM methods. Incorporating the unbiased and less noisy $U^{1}$ helps enhance performance in these longer horizons. This observation suggests that it is especially beneficial to use 2-UCBs in longer horizons.
@@
In summary, we find that LOLA outperforms existing approaches for experimentation by leveraging the strengths of LLMs and combining them with the benefits of active/online learning. In particular, our results demonstrate the value of LOLA as the time horizon increases, highlighting the limitations of relying solely on pure LLM predictions or the naive E&C approach.